# Small Drafts, Big Verdict: Information-Intensive Visual Reasoning via Speculation

### v1

@@ E.1 Impact of visual input to verdict
Method
InfographicVQA
ANLS
ChartMuseum
Acc
ChartQAPro
Acc
@@
GPT-4o Verdict w input 88.4 49.3 64.0
E.2 Impact of structured image input to verdict
In our experimental setup in Section 4.1, we preprocess each image via PP-StructureV3, a document parsing model that generates Markdown representations capturing layout, textual blocks, and visual metadata (Cui et al., 2025). This structured representation is then rendered as an image and provided as an additional image input for the verdict. This allows the verdict to access both the raw visual content and a layout-aware text representation simultaneously. To verify whether this input is critical or merely auxiliary, we conduct an ablation study (Table 10).
@@ E.2 Impact of structured image input to verdict
GPT-4o Verdict w input 88.4 49.3 64.0
Appendix F Error Analysis of Tool-driven Pipeline
@@
(iii) Lack of robustness on long and dense images. Information-intensive images often contain multi-panel figures and dense annotations. DeepEyes cannot maintain a trajectory across multiple zoom steps, making it difficult to integrate dispersed evidence. As a result, tasks requiring cross-region synthesis, such as counting, sorting, or comparing across multiple subplots, remain challenging for it.
@@
Appendix G Qualitative example
@@
Appendix H Prompt templates
H.1 Chain-of-Thought Prompts
As described in Section 4.1, we employ a Chain-of-Thought prompt for each consensus expert to generate reasoning paths and apply it identically when evaluating baselines. For InfographicVQA and HR-Bench 4K, we use the same CoT prompt. For ChartMuseum (Tang et al., 2025), we adopt its official reasoning prompt, and adapt that prompt strategy to ChartQAPro, given their similarity in task complexity. Since ChartQAPro requires different prompt templates tailored to question types (Masry et al., 2025), we first follow its official template per question type, then concatenate it with our reasoning prompt.
@@ H.2 Prompts for Verdict
Appendix I The Use of Large Language Models (LLMs)
In this work, we used LLMs solely for auxiliary tasks such as language polishing, prompt refining, and proofreading. Importantly, these interventions did not contribute any main scientific insight, experimental design, or methodological advance. All core ideas, experiments, analyses, and claims in this paper are the work of the authors.

### v2

@@ H.1 Impact of Visual Input to Verdict
Method InfographicVQA ANLS ChartMuseum Acc ChartQAPro Acc
@@
GPT-4o Verdict w/ input (SV) 88.4 49.3 64.0
@@
GPT-4o Verdict w input (SV) 88.4 49.3 64.0
H.2 Impact of Structured Image Input to Verdict
In our experimental setup in Section 4.1, we preprocess each image via PP-StructureV3, a document parsing model that generates Markdown representations capturing layout, textual blocks, and visual metadata (cui2025paddleocr). This structured representation is then rendered as an image and provided as an additional image input for the verdict. This allows the verdict to access both the raw visual content and a layout-aware text representation simultaneously. To verify whether this input is critical or merely auxiliary, we conduct an ablation study (Table 21).
Appendix I Quantitative Analysis of Draft Complementarity
To characterize the complementarity of draft experts, we focus on minority-correct recovery cases on InfographicVQA where only one of the three selected experts is correct and SV subsequently recovers the correct answer. These cases are most diagnostic for understanding how specific models provide unique, correct information and how others behave to help the verdict distinguish cues.
We randomly sample 50 minority-correct recovery instances and manually categorize their dominant reasoning bottlenecks into five types, summarized in Table 22. Extraction-related failures account for half of the cases, followed by color matching and global scan.
Table 22: Distribution of dominant reasoning bottlenecks over 50 minority-correct recovery cases on InfographicVQA.
Reasoning bottleneck Description Frequency (%)
Extraction Locating and reading fine-grained text/numbers 50
Color matching Matching colors in legends/charts to labels 18
Global scan Aggregating evidence across the full image 16
Localization Finding query-relevant regions 10
Numerical comparison Comparing numerical values 4
Table 23 reports per-model success rates within each bottleneck category over the 50 minority-correct instances. We observe a clear division of labor across experts: GLM-4.1V-Thinking and MiMO-VL-RL are most reliable for fine-grained extraction (58% and 70% success, respectively), with MiMO additionally strong on global-scan cases (75%). This advantage is consistent with their reasoning-oriented behavior: their step-by-step trajectories enable iterative verification of extracted values and cross-checking across multiple regions. Ovis2.5 and GLM-4.1V-Thinking perform best on color matching (57% and 60%), while Qwen2.5-VL dominates localization (80%) and numerical comparison (100%). We find that other experts often fail in these categories due to keyword-comprehension errors, whereas Qwen2.5-VL’s correct grounding makes its answer salient for the verdict to identify even when other models provide detailed but misdirected reasoning. Overall, no single expert dominates all reasoning types, confirming that SV benefits from specialized and complementary strengths rather than redundant voting.
Table 23: Per-model success rates (%) on minority-correct recovery cases, broken down by reasoning bottleneck.
Reasoning type Qwen2.5-VL-Instruct GLM-4.1V-Thinking MiMO-VL-RL Ovis2.5 InternVL3
Extraction 15 58 70 38 15
Color matching 0 60 33 57 25
Localization 80 0 0 33 0
Global scan 0 33 75 20 0
Numerical comparison 100 0 0 0 0
Finally, we summarize the most frequent expert combinations all 243 minority-correct recovery cases on InfographicVQA, as shown in Table 24. Two typical complementarity patterns emerge. First, balanced visual-skill combinations (e.g., Qwen–Ovis–InternVL) appear in cases requiring diverse perceptual cues. Second, extraction-focused pairings (e.g., Qwen–MiMO–InternVL and Qwen–GLM–Ovis) arise in extraction-intensive problems, where reasoning drafts provide accurate fine-grained values and non-reasoning drafts contribute robust localization for cross-verification. Overall, this analysis supports our claim that SV does not simply average similar models: it exploits complementary reasoning strengths across draft experts, and the minority-correct recoveries we observe are closely associated with diverse and specialized reasoning trajectories across models.
Table 24: Most frequent expert combinations among all minority-correct cases on InfographicVQA.
Rank Model combination Frequency (%)
1 Qwen + Ovis + InternVL 33.3
2 Qwen + MiMO + InternVL 14.0
3 Qwen + GLM + Ovis 11.1
Appendix J Failure Mode Analysis of SV
We analyze cases where the majority of selected draft experts are correct but SV produces an incorrect answer on InfographicVQA with GPT-4o as verdict. These instances isolate failures when the consensus mechanism or verdict synthesis fails despite having correct information available. Under exact-match scoring, such cases account for 3.98% of majority-correct examples; after normalizing answer formats (e.g., “Feb 11” vs. “February 11”), the true failure rate is 1.34% (34 cases).
Failure modes. Two dominant patterns emerge. (i) Fine-grained extraction errors (52.9%) arise from tiny numbers or small text on very tall infographics. Although draft experts often extract the correct cue, GPT-4o may fail to verify it because its tile-based high-resolution pipeline downsamples long inputs, losing subtle details. (ii) Color matching errors (35.3%) require aligning colored legends or regions with labels. This reflects a shared VLM capability gap: when both drafts and verdict struggle with color discrimination, drafts provide uncertain reasoning (sometimes explicit guesses), and GPT-4o exhibits verbosity bias in 58.3% of such cases (i.e., it follows the draft with the longest trajectory despite majority consensus pointing elsewhere).
Overall, SV inherits current VLM weaknesses in precise color reasoning and fine-grained perception, and when visual evidence is ambiguous the verdict may over-weight fluent yet incorrect drafts.
Table 25: Breakdown of SV failure modes on InfographicVQA.
Failure mode Proportion
Fine-grained extraction 52.9%
Color matching 35.3%
Other 11.8%
Appendix K Failure Mode Analysis of Tool-Driven Pipeline
@@ Appendix I Quantitative Analysis of Draft Complementarity
(iii) Lack of robustness on long and dense images. Information-intensive images often contain multi-panel figures and dense annotations. DeepEyes cannot maintain a trajectory across multiple zoom steps, making it difficult to integrate dispersed evidence. As a result, tasks requiring cross-region synthesis, such as counting, sorting, or comparing across multiple subplots, remain challenging.
@@
Appendix L Qualitative Example
@@
Appendix M Prompt Templates
M.1 Chain-of-Thought Prompts
As described in Section 4.1, we employ a Chain-of-Thought prompt for each consensus expert to generate reasoning paths and apply it identically when evaluating baselines. For InfographicVQA and HR-Bench 4K, we use the same CoT prompt. For ChartMuseum (tang2025chartmuseum), we adopt its official reasoning prompt, and adapt that prompt strategy to ChartQAPro, given their similarity in task complexity. Since ChartQAPro requires different prompt templates tailored to question types (masry2025chartqapro), we first follow its official template per question type, then concatenate it with our reasoning prompt.
@@
Appendix N The Use of Large Language Models (LLMs)
In this work, we used LLMs solely for auxiliary tasks such as language polishing, prompt refining, and proofreading. Importantly, these interventions did not contribute any main scientific insight, experimental design, or methodological advance. All core ideas, theoretical developments, experiments, analyses, and claims in this paper are the work of the authors. Accordingly, authors assume full responsibility for all content, consistent with ICLR policy that LLMs are not eligible for authorship.